OpenStudy (anonymous):

here are 3 questions i also need help with

OpenStudy (angela210793):

where?

OpenStudy (anonymous):

here

OpenStudy (angela210793):

(8x-2)/(x^2-1)=7x/(x^2-1) 8x-2=7x x=2 (a-9)/(a+4)=(a^2-3)/(a+4) a-a^2-6=0 this has solution only in complex no.s (2a-1)=(a+7)/(a+1) cross multiply 2a^2+a-1-a-7=0 2a^2-8=0 a^2-4=0 (a-2)(a+2)=0 so a=2 or a=-2